OpenStudy (anonymous):

A die is rolled three times. Find the chance that three different faces appear.

OpenStudy (anonymous):

whats the probability of one face?

OpenStudy (anonymous):

1*5/6*4/6=20/36=10/18=5/9 chance i think that's the answer

OpenStudy (anonymous):

@Khushal_Shah let him get to the result. Don't give it at the beginning

OpenStudy (anonymous):

that's right thanks :)

OpenStudy (anonymous):

@Khushal_Shah i wanna ask why is 5 and 4 ?

OpenStudy (anonymous):

@gozde35 there are 6 sides on a dice, right? So after the first roll, how many other sides are left to roll that you haven't seen yet?

OpenStudy (anonymous):

@Khushal_Shah hmm i see now 6/6 * 5/6 * 4/6 then it's 5/9 thank you :)

OpenStudy (anonymous):

yep, anytime man :)